A basketball leaves a player's hands at a height of 2.10 m above the floor. The basket is 3.05 mabove the floor. The player likes to shoot the ball at a 40.0 ∘ angle.

If the shot is made from a horizontal distance of 6.60 m and must be accurate to ±0.23m (horizontally), what is the range of initial speeds allowed to make the basket?
Answer #1

Case 1 :

X = horizontal distance

vox = component of launch velocity along x-direction = vo Cos40

t = time of travel

Time of travel is given as

t = X/vox (Since there is no acceleration along x-direction)

t = X/(vo Cos40)

Consider the motion along the Y-direction :

Yo = initial position = 2.10 m

Yf = final position = 3.05 m

a = acceleration = - 9.8 m/s2

voy = component of launch velocity along x-direction = vo Sin40

t = time of travel = X/(vo Cos40)

using the equation

Yf = Yo + voy t + (0.5) a t2

3.05 = 2.10 + (vo Sin40) (X/(vo Cos40)) + (0.5) (- 9.8) (X/(vo Cos40))2

3.05 = 2.10 + (Sin40) (X/(Cos40)) + (0.5) (- 9.8) (X/(vo Cos40))2

when X = 6.60 - 0.23 = 6.37 m

3.05 = 2.10 + (Sin40) (6.37/(Cos40)) + (0.5) (- 9.8) (6.37/(vo Cos40))2

vo = 8.8 m/s

when X = 6.60 + 0.23 = 6.83 m

3.05 = 2.10 + (Sin40) (6.83/(Cos40)) + (0.5) (- 9.8) (6.37/(vo Cos40))2

vo = 8.42 m/s

so the range is from 8.42 m/s to 8.8 m/s